Circular No. 5614

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CYGNI 1992 S. N. Shore and T. B. Ake, III, Goddard High Resolution Spectrograph

(GHRS) team and Computer Sciences Corporation (CSC); S. Starrfield, Arizona State University; G. Sonneborn, Goddard Space Flight Center, NASA; R. Gonzalez-Riestra, International Ultraviolet Explorer Observatory, European Space Agency, Madrid; S. Kraemer, Space Telescope Imaging Spectrograph team and CSC; and B. Altner, Applied Research Corporation report: “Observations of Nova Cyg 1992 were obtained with the GHRS on Sept. 8.2 UT with a spectral resolution of 15 000. The line profiles of N IV] 148.7-nm, Ne IV] 160.2-nm, He II 164.0-nm, N III] 175.0-nm, Si III] 189.2-nm, and C III] 190.8-nm show that the nova ejecta are composed of a complex of individual knots. These are nearly symmetric about line center. The nonresonance lines all have a saddle appearance, which suggests that the ejecta are nonspherical and probably bipolar. The individual emission features are resolved with velocity widths ranging between 100 and 200 km/s. Ionization variations are detected using the N IV] to N III] ratio. The highest velocity ejecta are more ionized than those at low velocity. The Si III] to C III] ratio indicates that the approaching material has a higher density than the receding part of the ejecta. He II 164.0-nm is broader than any of the other lines, with a velocity width of ∼ 3700 km/s (full width zero intensity). He II and Ne IV] show similar line profile features, suggesting that the ejecta are well mixed in neon. The GHRS spectra show a flat continuum at ∼ 5 × 10^{−13} erg cm^{−2} s^{−1} A^{−1}. Contemporaneous observations taken on Sept. 7.7 with the International Ultraviolet Explorer satellite yield the following fluxes (10^{−11} erg cm^{−2} s^{−1}): Lyα 121.6-nm, 2.40; N V 124.0-nm, 9.58; Si IV/O IV 140.0-nm, 8.18; N IV] 148.7-nm, 17.39; C IV 155.0-nm, 5.75; Ne V] 157.5-nm, 1.64; Ne IV] 160.2-nm, 6.28; He II 164.0nm, 4.90; O III] 166.5-nm, 8.79; N III] 175.0-nm, 14.42; Si III] 189.2-nm, 0.31; and C III] 190.8-nm, 3.17. Al III 186.0-nm is definitely present. The integrated 120- to 200-nm flux continues to decline at a steady rate of ∼ 10 percent per week. Long-wavelength observations show the presence of a broad [Na V] 207.0-nm line. There is a strong unidentified line at 215.0 nm that has been present since the ejecta became optically thin.”

Selected visual magnitude estimates: Sept. 6.78 UT, 9.4 (E. Schweitzer, Strasbourg, France); 8.16, 9.0 (W. G. Dillon, Missouri City, TX); 11.99, 9.0 (B. H. Granslo, Fjellhamar, Norway).

1992 September 17 Daniel W. E. Green